# COMPASS: Computational Mapping of Patient-Therapist Alliance Strategies with Language Modeling

### v2

@@ Abstract
The therapeutic working alliance is a critical factor in predicting the success of psychotherapy treatment. Traditionally, working alliance assessment relies on questionnaires completed by both therapists and patients. In this paper, we present COMPASS, a novel framework to directly infer the therapeutic working alliance from the natural language used in psychotherapy sessions. Our approach utilizes advanced large language models (LLMs) to analyze transcripts of psychotherapy sessions and compare them with distributed representations of statements in the working alliance inventory. Analyzing a dataset of over 950 sessions covering diverse psychiatric conditions including anxiety, depression, schizophrenia, and suicidal tendencies, we demonstrate the effectiveness of our method in providing fine-grained mapping of patient-therapist alignment trajectories and offering interpretability for clinical psychiatry and in identifying emerging patterns related to the condition being treated. By employing various deep learning-based topic modeling techniques in combination with prompting generative language models, we analyze the topical characteristics of different psychiatric conditions and their evolution at a turn-level resolution. This combined framework enhances the understanding of therapeutic interactions, enabling timely feedback for therapists regarding the quality of therapeutic relationships and providing interpretable insights to improve the effectiveness of psychotherapy.
@@ 1 Introduction
The working alliance, which encompasses various cognitive and emotional aspects of the therapist-patient relationship, is a critical concept in psychotherapy identified as a crucial factor in predicting treatment outcomes [1, 2]. Research over the past few decades has consistently demonstrated the significant role that a strong working alliance plays in determining the success of various therapeutic modalities, especially in patient engagement and adherence to treatment [3, 4, 5]. However, current methods for assessing the alliance rely on evaluating entire therapy sessions using point-scale ratings [6], with questionnaires and clinical instruments such as the working alliance inventory (WAI). These methods can be time-consuming and lack scalability for analyzing individual dialogue turns. Psychotherapy was among the first disciplines to embrace Natural Language Processing (NLP), with early applications including chatbots like ELIZA and Parry, which simulated psychotherapists and patients with schizophrenia, respectively [7, 8]. However, none of these attempts have been adopted or even systematically researched.
As a current standard in the field, the working alliance inventory (WAI) is a self-report measurement designed to quantify the therapeutic bond, task agreement, and goal agreement in psychotherapy [6, 9, 10]. It is typically completed by both patient and therapist at multiple points during therapy to monitor the development and quality of the therapeutic alliance. Common practice involves administering the WAI at the beginning of treatment to establish a baseline, at mid-treatment to track progress, and toward the end to assess changes in the alliance over time. In some settings, it may also be completed periodically throughout therapy, such as after every few sessions or even after each session, to continuously evaluate alliance strength.
@@
Our study leverages the WAI as a foundation for analyzing patient-therapist interactions, specifically examining the trajectory of alliance scores over time across different psychiatric conditions. By applying automated analysis methods to infer alliance scores at a granular, turn-by-turn level, we extend the traditional use of WAI. This approach enables us to explore fine-grained dynamics that are typically overlooked in standard WAI applications. Such an analysis can provide actionable insights into how alliance trajectories may vary by diagnosis, offering a deeper understanding of how therapeutic bonds develop within sessions. Ultimately, this could inform tailored therapeutic strategies that align more closely with the unique needs of each patient group, enhancing the utility of WAI in diverse clinical contexts.
@@
Despite its utility, the WAI is limited by its scope and intermittent administration, which may miss within-session processes that are central to effective therapy, such as the moment-to-moment shifts in alliance that drive therapeutic progress. Traditional tools like the WAI provide valuable insights but often lack the granularity to capture these intricate dynamics in real-time. To address this gap, we propose here an approach for quantifying patient-therapist alliance by leveraging large language models (LLMs) to project each dialogue turn onto representations of established working alliance inventories [6, 9, 10]. Following recent advancements in NLP [20, 21, 22, 23, 24], our approach enables us to not only estimate the overall degree of alliance but also identify fine-grained patterns and dynamics across shorter and longer time scales, e.g., turn-by-turn in a session and across sessions.
We employ the Counseling and Psychotherapy Transcripts from the Alexander Street dataset [25], which includes transcribed recordings of over 950 therapy sessions between multiple anonymized therapists and patients with clinical diagnoses of anxiety, depression, schizophrenia, or suicidal tendencies. We introduce the Working Alliance Transformer (WAT) and the Working Alliance LSTM (WA-LSTM), a Transformer-based and a Long-Short Term Memory-based model, respectively, that leverage the Working Alliance Inventory (WAI) [6, 22], and demonstrate how our method identifies patterns in therapy transcripts discriminative of the different psychiatric conditions. Previous studies have demonstrated that NLP techniques, including topic modeling, can reveal latent structures within depression-related language collected from platforms like Twitter [26] and improve the detection of Post-Traumatic Stress Disorder [27]; to facilitate interpretable insights, we conduct a systematic investigation of the most prominent topics that are addressed at the turn level (Figure 1).
In summary, this article presents COMPASS (COmputational Mapping of Patient-Therapist Alliance StrategieS), a novel approach for estimating the therapeutic working alliance from regular sessions transcripts, using the WAI as a matching template. The approach is validated by improved classification and diagnostic capabilities of deep learning models over agnostic models. Our approach also offers interpretable insights that can inform psychotherapy strategies. By leveraging LLMs, we enable a more granular and comprehensive understanding of the working alliance, allowing for timely feedback and improved effectiveness in psychotherapy treatments.
@@ 2.2 Deep Learning Inference of Working Alliance
The analytic framework for inferring the working alliance from psychotherapy sessions is illustrated in Figure 1. Our approach involves processing the full records of individual patients with multiple clinical conditions or a cohort of patients with the same condition, which can be segmented based on timestamps or topic turns. The original data is presented in pairs of dialogues, and we extract features in three different ways: (1) using the full pairs of dialogues, (2) extracting only the patients’ responses, or (3) extracting only the therapists’ responses. Each feature set has its advantages and disadvantages. The dialogue features contain all the information but can mix together the intents within sentences from both individuals.
@@ 2.3 Sentence Embeddings
Table 1: Examples of working alliance inference
@@
2.4 Working Alliance Inventories
The Working Alliance Inventory (WAI) is a self-report measurement questionnaire designed to quantify the therapeutic bond, task agreement, and goal agreement in psychotherapy [6, 9, 10], typically completed by patient and therapist at multiple points during therapy to monitor the development and quality of the therapeutic alliance. Common practice involves completing it at the beginning of treatment to establish a baseline, mid-treatment to track progress, and toward the end to assess changes in the alliance over time. In some cases, it may also be completed periodically throughout therapy, such as after every few sessions, or even after each session to continuously evaluate the strength of the alliance.
It has been widely used to assess the quality of the working alliance between therapists and patients and has demonstrated good psychometric properties [9, 10]. Research indicates that higher WAI scores are generally associated with better therapy outcomes across various diagnoses, including depression and anxiety disorders. Effect sizes for the correlation between alliance quality and treatment outcomes range from 0.19 to 0.32, suggesting a moderate relationship [11, 12]. The effectiveness of the therapeutic alliance can vary depending on the type of therapy employed. For example, cognitive-behavioral therapy (CBT) often shows different patterns of alliance development compared to psychodynamic approaches. Studies suggest that in CBT, a strong initial alliance may be crucial for engagement and adherence to treatment [13]. The WAI has demonstrated high internal consistency (Cronbach’s alpha around 0.93) and has been validated across diverse populations and settings [11, 13].
The modern version of the WAI consists of 36 questions, and participants are asked to rate each item on a 7-point scale (1=never, 7=always) [10], as detailed by the full inventory in the supplementary material (Table S1). The working alliance inventories are usually administered with two versions, a client / patient version, and a therapist version. The inventory items are very similar with only a minor change of phrasing. The inventory aims to measure alliance factors across different types of therapy, establish the relationship between the alliance measure and theoretical constructs underlying the measure, and relate the alliance measure to a unified theory of therapeutic change [3]. Both patients and therapists fill out the WAI independently after each session, and responses are blinded to the other party to ensure unbiased assessments.
We then compute the sentence similarity (e.g., cosine similarity) of the 300-dimension embedding vector of the example sentence with the 300-dimension embedding vector of the inventory item, yielding a 36-dimensional similarity score. For example, the first sentence corresponds to [0.52, 0.49, 0.47, …], with 0.52 being the similarity between the first sentence and the first inventory item. Using the key table corresponding to the working alliance scores, we map the 36 items to the 3 scales, which gives the inferred scores as shown in Table 1. As expected, the first sentence shows a high score for BOND, the second for TASK and GOAL, and the third for GOAL, demonstrating the model’s ability to differentiate and quantify the various aspects of the therapeutic working alliance.
@@ 3.4 Major Themes in Psychotherapeutic Topics and Interventions
The result of this analysis is presented in Table 3, which lists the high-level description of each topic, the likely interventional rationale and a literal example. We found that the vast implications of these topics and interventions can be partially summarized four major themes, as follows:
@@ 3.5 Principal Topic Space in Psychotherapy Sessions
Variability Among Different Clinical Conditions. It is important to note that the four datasets exhibit variability in terms of the specific patient populations represented.
@@ 4.1 Working Alliance Analysis
While our approach proves effective, there are limitations inherent in inferring psychological states from text data. One potential limitation is the reliance on semantic similarity measures, which may not fully capture the directionality of similarity. For example, it may not differentiate between a statement that is irrelevant to the inventory item and a statement that is opposite to the inventory item. To address this, alternative approaches can incorporate counter arguments and compute similarity scores based on the difference between positive and negative similarity values. However, our findings suggest that the sentence embedding we use already captures the concept of negation, and we observe no clear difference in performance between the alternative approach and our prototypical approach.
@@ 4.2 Topic Modeling and Interpretability
However, it is important to note that the learned topics may exhibit variability across different datasets due to the specific patient populations represented. Each psychiatric condition has its unique challenges and therapeutic goals, which require tailored approaches and interventions by the therapists. Therefore, the topics identified within each condition reflect the prevalent themes and concerns specific to that condition.
@@ 5 Conclusions
In this study, we have introduced an approach that combines state-of-the-art language modeling with therapy-evaluation inventories to provide a detailed representation of the interaction between patients and therapists. Our method offers granular insights for post-session interpretations and has the potential to assist in diagnosing patients based on linguistic features. While our focus has been on the Working Alliance Inventory, our approach is generic and can be extended to other assessment instruments in the field of psychotherapy.
Additionally, we have made contributions in the area of deep learning-based topic modeling to further enhance our analysis. Our first objective was to compare various neural topic modeling methods in learning the topical propensities of different psychiatric conditions. We found that different coherence measures yield different rankings of the topic models, but there are a few models, such as Wasserstein Topic Models and Embedded Topic Models, that perform well in terms of coherence and diversity.
Furthermore, we have incorporated temporal modeling into topic modeling to parse topics in different segments of the therapy sessions. This temporal analysis adds another layer of interpretability and enables us to observe session trajectories and their separability between patients and therapists. We have noted that in sessions with patients with anxiety and depression, the trajectories of patients and therapists tend to be more separable, whereas in schizophrenia sessions, they are more entangled. This initial step toward turn-level resolution temporal analysis in topic modeling provides valuable insights that can help therapists improve the effectiveness of psychotherapy.
In conclusion, our combined framework of working alliance analysis and topic modeling offers interpretable insights for therapists. By leveraging language models and incorporating temporal analysis, we aim to enhance the understanding of the therapeutic process and support therapists in providing more effective and personalized care to their patients.
@@ Acknowledgments
The authors thank Drs. Barnaby Nelson, Alison Yung, and Ravi Tejwani for helpful discussions and suggestions.

### v3

@@ Abstract
The therapeutic working alliance is a critical predictor of psychotherapy success. Traditionally, working alliance assessment relies on questionnaires completed by both therapists and patients. In this paper, we present COMPASS, a novel framework to directly infer the therapeutic working alliance from the natural language used in psychotherapy sessions. Our approach leverages advanced large language models (LLMs) to analyze session transcripts and map them to distributed representations. These representations capture the semantic similarities between the dialogues and psychometric instruments, such as the Working Alliance Inventory. Analyzing a dataset of over 950 sessions spanning diverse psychiatric conditions – including anxiety (N=498), depression (N=377), schizophrenia (N=71), and suicidal tendencies (N=12) – collected between 1970 and 2012, we demonstrate the effectiveness of our method in providing fine-grained mapping of patient-therapist alignment trajectories, offering interpretable insights for clinical practice, and identifying emerging patterns related to the condition being treated. By employing various deep learning-based topic modeling techniques in combination with prompting generative language models, we analyze the topical characteristics of different psychiatric conditions and how these topics evolve during each turn of the conversation. This integrated framework enhances the understanding of therapeutic interactions, enables timely feedback for therapists on the quality of therapeutic relationships, and provides clear, actionable insights to improve the effectiveness of psychotherapy.
@@ 1 Introduction
The working alliance, which encompasses various cognitive and emotional aspects of the therapist-patient relationship, is a critical concept in psychotherapy identified as a crucial factor in predicting treatment outcomes [1, 2]. Research over the past few decades has consistently demonstrated the significant role that a strong working alliance plays in determining the success of various therapeutic modalities, especially in patient engagement and adherence to treatment [3, 4, 5]. However, current methods for assessing the alliance rely on evaluating entire therapy sessions using point-scale ratings [6], with questionnaires and clinical instruments such as the working alliance inventory (WAI). As a current standard in the field, the working alliance inventory (WAI) is a self-report measurement designed to quantify the therapeutic bond, task agreement, and goal agreement in psychotherapy [6, 7, 8]. It is typically completed by both patient and therapist at multiple points during therapy to monitor the development and quality of the therapeutic alliance. Common practice involves administering the WAI at the beginning of treatment to establish a baseline, at mid-treatment to track progress, and toward the end to assess changes in the alliance over time. In some settings, it may also be completed periodically throughout therapy, such as after every few sessions or even after each session, to continuously evaluate alliance strength.
@@
While the WAI is a valuable tool, its periodic administration cannot capture the nuanced, moment-to-moment alliance dynamics that drive therapeutic progress. Furthermore, regular WAI assessment proves too time-intensive for both clinical practice and large-scale research. Although psychotherapy was an early adopter of Natural Language Processing (NLP), with pioneering efforts like ELIZA and Parry simulating therapeutic interactions [18, 19], these innovations never gained systematic clinical adoption.
We propose leveraging large language models (LLMs) to quantify patient-therapist alliance by mapping dialogue turns onto established WAI dimensions [6, 7, 8]. Building on recent NLP advances [20, 21, 22, 23, 24], our approach enables analysis of alliance patterns across multiple timescales - from individual turns to entire therapeutic trajectories. Our methodology, COMPASS (COmputational Mapping of Patient-Therapist Alliance StrategieS), analyzes therapy transcripts by mapping each conversational turn to WAI concepts, generating granular alliance scores that can reveal diagnosis-specific patterns in therapeutic bond development [22]. We validate this approach using the Alexander Street dataset [25], comprising over 950 transcribed sessions with patients diagnosed with anxiety (N=498), depression (N=377), schizophrenia (N=71), and suicidal tendencies (N=12). Previous research has demonstrated NLP’s utility in analyzing mental health discourse, from uncovering latent structures in depression-related language on Twitter [26] to enhancing PTSD detection [27].
Building on this foundation, we conduct systematic topic modeling at the dialogue turn level to generate interpretable insights (Figure 1). COMPASS’s effectiveness is validated through improved classification and diagnostic capabilities compared to baseline models, offering actionable insights for enhancing psychotherapy strategies through timely, granular alliance monitoring.
@@ 2.2 Deep Learning Inference of Working Alliance
The analytic framework for inferring the working alliance from psychotherapy sessions is illustrated in Figure 1. Our approach involves analyzing the complete set of transcript records, which can be segmented based on timestamps or topic turns, from individual patients with various clinical conditions or from cohorts of patients with the same condition. The original data is presented in pairs of dialogues, and we extract features in three different ways: (1) using the full pairs of dialogues, (2) extracting only the patients’ responses, or (3) extracting only the therapists’ responses. Each feature set has its advantages and disadvantages. The dialogue features contain all the information but can mix together the intents within sentences from both individuals.
@@ 2.3 Sentence Embeddings
2.4 Working Alliance Inventory
The modern version of the working alliance inventory (WAI) consists of 36 questions, and participants are asked to rate each item on a 7-point scale (1=never, 7=always) [8], as detailed by the full inventory in the supplementary material (Table S1). The working alliance inventories are usually administered with two versions, a client / patient version, and a therapist version. The inventory items are very similar with only a minor change of phrasing. The inventory aims to measure alliance factors across different types of therapy, establish the relationship between the alliance measure and theoretical constructs underlying the measure, and relate the alliance measure to a unified theory of therapeutic change [3]. Both patients and therapists fill out the WAI independently after each session, and responses are blinded to the other party to ensure unbiased assessments.
We then compute the sentence similarity (e.g., cosine similarity) of the 300-dimension embedding vector of the example sentence with the 300-dimension embedding vector of the inventory item, yielding a 36-dimensional similarity score. For example, the first sentence corresponds to [0.52, 0.49, 0.47, …], with 0.52 being the similarity between the first sentence and the first inventory item. Using the key table corresponding to the working alliance scores, we map the 36 items to the 3 scales by calculating the dot product between the 36-dimensional similarity vector and the corresponding key table vector (e.g., [-1, 1, 0, 1, +1, 0, …]). That is, the semantic similarity of the turns with respect to each item is summed up following the original aggregation of the WAI score (Table S1). This mapping produces the inferred scores for Task, Bond, and Goal, as presented in Table 1.
Table 1: Examples of working alliance inference. The semantic similarities of the sentences to the scales’ items are summed up following the original WAI score aggregation (S1).
@@ 2.4 Working Alliance Inventory
As expected, the first sentence shows a high score for BOND, the second for TASK and GOAL, and the third for GOAL, demonstrating the model’s ability to differentiate and quantify the various aspects of the therapeutic working alliance.
@@ 3.4 Major Themes in Psychotherapeutic Topics and Interventions
The result of this analysis is presented in Table 3, which lists the high-level description of each topic, the likely interventional rationale and a literal example. We found that the vast implications of these topics and interventions can be partially summarized into four major themes with the prompt “can you summarize them into a few major themes?”, as follows:
@@ 3.5 Principal Topic Space in Psychotherapy Sessions
Variability Among Different Clinical Conditions. Our dataset comprises four subsets, each representing a distinct patient population characterized by a specific clinical condition. To capture this variability, we trained individual topic models for each subpopulation cohort. The learned topics were analyzed to highlight condition-specific themes and insights.
@@ 4.1 Working Alliance Analysis
Limitations. While our approach proves effective, there are limitations inherent in inferring psychological states from text data. One potential limitation is the reliance on semantic similarity measures, which may not fully capture the directionality of similarity. For example, it may not differentiate between a statement that is irrelevant to the inventory item and a statement that is opposite to the inventory item. To address this, alternative approaches can incorporate counter arguments and compute similarity scores based on the difference between positive and negative similarity values. However, our findings suggest that the sentence embedding we use already captures the concept of negation, and we observe no clear difference in performance between the alternative approach and our prototypical approach.
@@ 4.2 Topic Modeling and Interpretability
Limitations. However, it is important to note that the learned topics may exhibit variability across different datasets due to the specific patient populations represented. Each psychiatric condition has its unique challenges and therapeutic goals, which require tailored approaches and interventions by the therapists. Therefore, the topics identified within each condition reflect the prevalent themes and concerns specific to that condition.
@@
A final limitation is the lack of validation of the interpretation results we obtained using LLMs. A proper validation by human experts is beyond the scope of the present study, and would require a significant effort of marshaling human resources and designing a systematic approach that can accommodate the expected increase in LLM capabilities. We hope that our study can contribute to initiate such effort.
@@ 5 Conclusions
In this study, we have introduced an approach that combines state-of-the-art language modeling with therapy-evaluation inventories to provide a detailed representation of the interaction between patients and therapists. Our method offers granular insights for post-session interpretations and has the potential to assist in diagnosing patients based on linguistic features. This dynamic analysis adds a layer of interpretability to the traditional session-based WAI evaluation, and enables us to observe session trajectories and their separability between patients and therapists. We have noted, for instance, that in sessions with patients with anxiety and depression, the trajectories of patients and therapists tend to be more separable, whereas in schizophrenia sessions, they are more entangled. This initial step toward turn-level resolution temporal analysis in topic modeling provides valuable insights that can help therapists improve the effectiveness of psychotherapy. Moreove, while our focus has been on the Working Alliance Inventory, our approach is generic and can be extended to other assessment instruments in the field of psychotherapy.
In conclusion, our analytic framework offers interpretable insights for therapists. By leveraging language models and incorporating temporal analysis, we aim to enhance the understanding of the therapeutic process and support therapists in providing more effective and personalized care to their patients.
@@ Acknowledgments
The authors thank Drs. Barnaby Nelson, Alison Yung, and Ravi Tejwani for their helpful discussions and suggestions. We also sincerely appreciate the anonymous reviewers and the editorial team for their constructive feedback and insightful suggestions, which have helped improve this manuscript. Additionally, we acknowledge our collaborators and colleagues for their valuable input and support throughout this study. BL is supported by the Hasso Plattner Foundation, the Windreich Family Foundation, and the National Institutes of Health (NIH). This work was supported in part by the NIH IMPACT-MH PREDiCTOR grant 1U01MH136535-01, which provided funding for BL, YL, RJ, CC, and GC.
@@
Conflict of Interest Statement
DB and GC are employees of IBM. All other authors declare no competing interests.
Data Availability Statement
The data analyzed in this study are publicly available from previously published sources. Further details and access information are provided in the references or available upon request.
Ethics Approval and Consent to Participate Statement
This study involved secondary analysis of de-identified, publicly available data. All methods were performed in accordance with relevant guidelines and regulations. As no new data collection from human participants was conducted by the authors, ethics approval and informed consent were not required. Ethical approval and participant consent were obtained by the original data providers, as detailed in the original publications.